OpenStudy (anonymous):

Solve the equation with the replacement set {–76, –8, 8, 76} |42 – x | = 34 A. 8 B. 76 C. -8, 8 D. 8, 76 please! thank you!

OpenStudy (nurali):

|42 − x| = 34 |x − 42| = 34 x − 42 = ± 34 x = 42 ± 34 x = 8, 76